# Can RLHF be More Efficient with Imperfect Reward Models? A Policy Coverage Perspective

### v2

@@ Abstract
Sample efficiency is critical for online Reinforcement Learning from Human Feedback (RLHF). While existing works investigate sample-efficient online exploration strategies, the potential of utilizing misspecified yet relevant reward models to accelerate learning remains underexplored. This paper studies how to transfer knowledge from those imperfect reward models in online RLHF. We start by identifying a novel property of the KL-regularized RLHF objective: a policy’s coverability of the optimal policy is captured by its sub-optimality. Building on this insight, we propose novel transfer learning principles and a theoretical algorithm with provable benefits compared to standard online learning. Our approach achieves low regret in the early stage by quickly adapting to the best available source reward models without prior knowledge of their quality, and over time, it attains an $\widetilde{O}(\sqrt{T})$ regret bound independent of structural complexity measures. Empirically, inspired by our theoretical findings, we develop a win-rate-based transfer policy selection method with improved computational efficiency. Moreover, our empirical transfer learning technique is modular and can be integrated with various policy optimization methods, such as DPO, IPO and XPO, to further enhance their performance. We validate the effectiveness of our method through experiments on summarization tasks.
@@ 1 Introduction
Motivated by these considerations, this paper studies how imperfect reward models can be used to learn a near-optimal policy with fewer human annotations. We consider the case where several source reward models are available, yet their quality, i.e., the similarity to human rewards, is unknown a priori. Our contributions are summarized as follows.
@@
  3. To reduce computational overheads and improve scalablity to practical RLHF scenarios, in Sec. 5, we further propose an empirical algorithm that selects transfer policy based on the win rates competing with the learning policy, rather than relying on the estimated policy values. On the one hand, win rates can be estimated much more efficiently than the policy values. On the other hand, as justified by our theory, win rates help to identify lower bounds for the coverability of the optimal policy. Notably, our empirical TPO is general: its core transfer learning technique can be modularized and combined with various reward-model-free policy optimization methods, such as DPO (Rafailov et al., 2024), IPO (Azar et al., 2024), XPO (Xie et al., 2024), to boost their performance. The effectiveness of our approach is demonstrated in fine-tuning T5 models on summarization tasks in Sec. 6.
@@ 1.1 Related Work
We only summarize the most relevant results on sample complexity and transfer RL here due to space limitation. Other closely related topics can be found in Appx. B.2.
@@ 3.2 New Insights for Transfer Learning in RLHF
To understand the significance of Thm. 3.2, consider the case when $\{\pi^{t}\}_{t\in[T]}$ are produced by a no-regret online learning algorithm, such as XPO in Lem. 2.4. As a result, the term $\sum_{t=1}^{T}\frac{J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{t})}{\beta T}$ in Eq. (4) diminishes to 0 as $T$ increases. This implies that the policy $\pi_{\text{Dstl}}$ distilled from online data by offline learning techniques converges to $\pi^{*}_{r^{*}}$ at a rate of $O(T^{-\frac{1}{2}})$, which does not depend on $|\mathcal{S}|,|\mathcal{A}|$ or other complexity measures such as ${\texttt{Cov}}_{\infty}(\Pi)$ in Lem. 2.4. This result not only strictly improves the sample complexity bounds [^3] Beyond sample complexity, a regret bound improved to $\widetilde{O}(\sqrt{T})$ for online RLHF can be established. We defer it to Coro. LABEL:coro:sqrtT_reg after presenting our main results. for existing online RLHF algorithms (Xiong et al., 2024; Xie et al., 2024; Cen et al., 2024; Zhang et al., 2024), but also reveals a fundamental difference from the pure reward maximization setting, where lower bounds depending on those structural complexity factors have been established (Auer et al., 2002; Dani et al., 2008). We defer detailed discussions to Appx. E.3.
@@ 4 Provably Efficient Transfer Learning
Outline of Main Algorithm  Our main algorithm TPO, short for Transfer Policy Optimization, is provided in Alg. 1, which leverages Alg. 2 (TPS, short for Transfer Policy Selection) as a subroutine to select source policies to transfer from. TPO can be regarded as a mixture of standard online learning and transfer learning, balanced through a hyper-parameter $\alpha\in(0,1)$. Motivated by the implication of Thm. 3.2, TPO returns the offline policy computed with all the data collected. For convenience, we divide the total number of iterations $T$ into $K=T/N$ blocks, each containing $N$ sub-iterations. In each block, we first run $\alpha N$ iterations of an OnLine learning algorithm $\texttt{Alg}_{\texttt{OL}}$, followed by $(1-\alpha)N$ iterations of transfer learning with policy selected by Alg. 2. Here, $\texttt{Alg}_{\texttt{OL}}$ can be any online algorithm with per-step no-regret guarantees, for example, XPO in Lem. 2.4. To save space, we defer to Appx. D the formal behavior assumption on $\texttt{Alg}_{\texttt{OL}}$ (Def. D.2) and concrete examples with verifications.
Now we are ready to take a closer look at the transfer policy selection steps in Alg. 2 in Sec. 4.1, after which, the provable merits of proposed approach are showcased in Sec. 4.2.
@@
             if $n\leq\alpha N$ then  $\pi^{k,n}\leftarrow\texttt{Alg}_{\texttt{OL}}(\alpha T,\Pi,\delta;\mathcal{D}^{k,n}_{\texttt{OL}})$ ;
             else   $\pi^{k,n}\leftarrow{{\texttt{TPS}}}(T,\Pi,\delta,\{r^{w}\}_{w=1}^{W};\mathcal{D}^{k,n})$ ;
@@
Before diving into details, we first clarify some notation. Given a dataset $\mathcal{D}:=\{(s^{i},a^{i},\widetilde{a}^{i},y^{i},\pi^{i})\}_{i=1}^{|\mathcal{D}|}$, $L_{\mathcal{D}}(r)$ denotes the average negative log-likelihood (NLL) loss regarding the reward model $r$:
@@
$\pi_{\text{Dstl}},{\widehat{r}}_{\text{Dstl}}\leftarrow{\texttt{RPO}}(\text{conv}(\Pi),\mathcal{R}^{\Pi},\mathcal{D},\eta)$ with $\eta=c\cdot(1+e^{{R}})^{-2}\sqrt{\frac{1}{|\mathcal{D}|}\log\frac{|\Pi|T}{\delta}}.$
@@
4.2 Main Results and Interpretation
@@
We decompose the total regret into two parts depending on their origins. $\text{Reg}_{\texttt{OL}}^{(t)}$ comes from the regret by running the online algorithm $\texttt{Alg}_{\texttt{OL}}$. It is weighted by $\alpha$ since we only allocate $\alpha$-proportion of the samples for $\texttt{Alg}_{\texttt{OL}}$. $\text{Reg}_{\texttt{Trf}}^{(t)}$ represents the regret from transfer policies. The first term in Eq. (8) reflects the benefits of utilizing transfer policy over online learning. Here $\Delta_{\min}$ is contributed by source reward models $\{r^{w}\}_{w\in[W]}$, and the term $\iota^{k(\tau),n(\tau)}$ is due to the “self-transfer policy” $\pi_{\text{Dstl}}$, as we derived in Lem. 4.2. The second term in Eq. (8) results from the imperfection of source reward models: without prior knowledge on their quality, additional cost has to be paid during exploration.
@@
Stage 1: $t<\frac{W^{2}}{\Delta^{2}_{\min}}$  This corresponds to the early learning stage, when $t$ is relatively small. In this case, Thm. 4.3 implies the following regret bound:
@@
Input: $K$, $N$ and $\{r^{w}\}_{w\in[W]}$;
@@
             $\forall w\in[W],~{}\widehat{\text{WR}}^{\pi^{*}_{r^{w}}}\leftarrow\widehat{{\mathbb{P}}}_{r^{*}}^{k,n}(\pi^{*}_{r^{w}}\succ\pi^{k}_{\texttt{OL}})+c\sqrt{\frac{1}{N^{k,n}(\pi^{*}_{r^{w}})}\log\frac{1}{\delta}}$;
@@
Choice of $\alpha$ and the total regret of TPO  Our algorithm treats $\alpha$ as a hyperparameter. Based on the discussion above, with proper choice of $\alpha$ we have the following corollary.
Corollary 4.4 (Total Regret of TPO).
By choosing XPO (Xie et al., 2024) as $\texttt{Alg}_{\texttt{OL}}$ and setting $\alpha=e^{-\frac{R}{\beta}}\leq{\texttt{Cov}}_{\infty}^{-1}(\Pi)$, Thm. 4.3 implies that TPO achieves $\widetilde{O}(W\sqrt{T})$ regret if $T<\frac{W^{2}}{\Delta^{2}_{\min}}$, or $\widetilde{O}(\sqrt{T})$ regret if $T>\frac{W^{2}}{\Delta^{2}_{\min}}$ and large enough.
Depending on the concrete scenarios, if we have stronger prior beliefs that one of $\{r^{w}\}_{w\in[W]}$ is similar to $r^{*}$, we should prefer larger $\alpha$. Besides, one may gradually decay $\alpha$ to 0 when the iteration number is large enough, which can also result in a total regret growing with $\widetilde{O}(\sqrt{T})$ over time.
Improved regret bound in standard online RLHF without source rewards Although our focus is transfer learning, our results can be extended to the standard online RLHF setting where no source tasks are present (i.e., $W=0$). As implied by Corollary 4.4, only utilizing self-transfer learning can result in an $\widetilde{O}(\sqrt{T})$ online regret, thereby strictly improving existing results (Xiong et al., 2024; Xie et al., 2024; Cen et al., 2024; Zhang et al., 2024).
@@ 5 From Theory to an Empirical Algorithm
Under BT-model444 Lem. 5.1 can be generalized beyond BT-model. Besides, it is possible to construct a lower bound involving ${\mathbb{P}}_{r^{*}}(\bar{\pi}\succ\pi)$ instead. See Lem. G.3 and Remark G.4 in Appx. G for more details. , for any $\pi$: ${\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi}\geq\!$$\max_{\gamma>0,\bar{\pi}}$$({\sqrt{(\gamma\!+\!2{\mathbb{P}}_{r^{*}}(\pi\!\succ\!\bar{\pi}))\log\frac{1+\gamma}{\gamma}}+\sqrt{\frac{J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\bar{\pi})}{2\beta}}})^{-1}.$
@@ 6 Experiments
Experiment Results and Interpretation  We run Alg. 3 for $K=3$ iterations and compare its performance with three baselines: (I) vanilla iterative-DPO without transfer learning (i.e., setting $\text{Alg}_{\text{PO}}$ = DPO and $W=0$); (II) purely exploiting the worst source reward—ROUGE score; (III) purely exploiting the best source reward—T5-large. Concretely, baseline (I) removes the transfer learning component in Alg. 3 by assigning $\pi^{k,n}=\pi_{\texttt{OL}}^{k}$ for all $n\in[N]$. For baselines (II) and (III), the worst (ROUGE-LSum) and best (T5-Large) reward models from the candidates (a)-(d) are selected, and pure transfer learning is then performed using the responses recommended by the chosen reward models, i.e., $\pi^{k,n}=\pi^{*}_{r^{w}}$ in Alg. 3 for the selected $r^{w}$. Here the worst and best reward models are selected based on the final policy value when aligning with the given reward model.
@@ Appendix C Offline Learning Results in Previous Literature
${\texttt{RPO}}(\widetilde{\Pi},\mathcal{R},\mathcal{D},\eta)=\arg\max_{\pi\in\widetilde{\Pi}}\min_{r\in\mathcal{R}}L_{\mathcal{D}}(r)+\eta\mathbb{E}_{s\sim\rho,a\sim\pi,\widetilde{a}\sim\pi_{\texttt{ref}}}[r(s,a)-r(s,\widetilde{a})]-\beta\text{KL}(\pi\|\pi_{\texttt{ref}}),$ (12)
@@ F.2 Some Useful Lemmas
F.3 Proof for Thm. 4.3
See 4.3 Throught the proof, we follow the convention that $1/0=+\infty$.
@@
We first bound the accumulative error when $\mathbb{I}[\mathcal{E}^{k,n}_{\text{Dstl}}]=1$. On the good events in Lem. 4.1 and Lem. 4.2 (which holds w.p. $1-\delta$), $\mathbb{I}[\mathcal{E}^{k,n}_{\text{Dstl}}]=1$ implies
@@
Here we only focus on those source tasks with $\Delta(w)>0$, since transferring from $\pi^{*}_{r^{w}}$ with $\Delta(w)=0$ does not incur regret. We separate source tasks into two sets $\mathcal{W}_{\leq 2\Delta_{\min}}:=\{w\in[W]|\Delta(w)\leq 2\Delta_{\min}\}$ and $\mathcal{W}_{>2\Delta_{\min}}:=\{w\in[W]|\Delta(w)>2\Delta_{\min}\}$. For $w\in\mathcal{W}_{>2\Delta_{\min}}$, on the same good events in Lem. 4.1 and Lem. 4.2, $\mathbb{I}[\mathcal{E}^{n,k}_{w}]=1$ implies
@@
On the other hand, on the good events of Lem. 4.1 and Lem. 4.2, when ${\mathbb{I}}[\mathcal{E}^{n,k}_{w}\cap\mathcal{E}^{n,k}_{2\iota<\Delta_{\min}}]=1$, we must have:
$J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{ref}})-\iota^{k,n}\leq$ ${\widehat{V}}^{k,n}(\pi^{k,n}_{\text{Dstl}};\mathcal{D}^{k,n-1}{})$ (Lem. 4.1 and Lem. 4.2)
@@
$\sum_{(k,n)\in\mathcal{I}_{\widetilde{K},\widetilde{N}}}J_{\beta}(\pi^{*}_{r^{*}})-$ $J_{\beta}(\pi^{k,n})=\widetilde{O}\Big{(}\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\Delta_{\min}\wedge\iota^{n,k}+{R}e^{2{R}}\sqrt{\alpha\mathcal{C}(\Pi)t_{\widetilde{K},\widetilde{N}}}+e^{2_{\max}}\sqrt{(1-\alpha)Wt_{\widetilde{K},\widetilde{N}}}\wedge\sum_{\begin{subarray}{c}w\in[W]\\ \Delta(w)>0\end{subarray}}\frac{e^{4{R}}}{\Delta(w)}\Big{)},$

### v3

@@ Abstract
Sample efficiency is critical for online Reinforcement Learning from Human Feedback (RLHF). While existing works investigate sample-efficient online exploration strategies, the potential of utilizing misspecified yet relevant reward models to accelerate learning remains underexplored. This paper studies how to transfer knowledge from those imperfect reward models in online RLHF. We start by identifying a novel property due to KL-regularization in the RLHF objective: a policy’s coverability of the optimal policy is captured by its sub-optimality. Building on this insight, we propose novel transfer learning principles and a theoretical algorithm—Transfer Policy Optimization (TPO)—with provable benefits compared to standard online learning. Empirically, inspired by our theoretical findings, we develop a win-rate-based transfer policy selection strategy with improved computational efficiency. Moreover, our empirical transfer learning technique is modular and can be integrated with various policy optimization methods, such as DPO, IPO and XPO, to further enhance their performance. We validate the effectiveness of our method through experiments on summarization tasks.
@@ 1 Introduction
Motivated by these considerations, this paper studies how to leverage imperfect reward models to learn a near-optimal policy with fewer human annotations. We consider the case where several source reward models are available, yet their quality, i.e., the similarity to human rewards, is unknown a priori. Our contributions are summarized as follows.
@@
  3. To reduce computational overheads and improve scalablity to practical RLHF scenarios, in Sec. 5, we propose an empirical version of TPO that selects transfer policy based on the win rates competing with the learning policy, rather than relying on the estimated policy values. On the one hand, win rates can be estimated much more efficiently than the policy values. On the other hand, as justified by our theory, win rates help to identify lower bounds for the coverability of the optimal policy. Notably, our empirical TPO is general: its core transfer learning technique can be modularized and combined with various reward-model-free policy optimization methods, such as DPO (Rafailov et al., 2024), IPO (Azar et al., 2024), XPO (Xie et al., 2024), to boost their performance. The effectiveness of our approach is demonstrated in fine-tuning T5 models on summarization tasks in Sec. 6.
@@ 1.1 Related Work
We summarize the most related ones on sample complexity and transfer RL here and defer the others to Appx. B.2.
@@ 3.2 New Insights for Transfer Learning in RLHF
To understand the significance of Thm. 3.2, consider the case when $\{\pi^{t}\}_{t\in[T]}$ are produced by a no-regret online learning algorithm, such as XPO in Lem. 2.4. As a result, the term $\sum_{t=1}^{T}\frac{J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{t})}{\beta T}$ in Eq. (4) diminishes to 0 as $T$ increases. This implies that the policy $\pi_{\text{Dstl}}$ distilled from online data by offline learning techniques converges to $\pi^{*}_{r^{*}}$ at a rate of $O(T^{-\frac{1}{2}})$, which does not depend on $|\mathcal{S}|,|\mathcal{A}|$ or other complexity measures such as ${\texttt{Cov}}_{\infty}(\Pi)$ in Lem. 2.4. This result not only strictly improves the sample complexity bounds [^3] Beyond sample complexity, a regret bound improved to $\widetilde{O}(\sqrt{T})$ for online RLHF can be established. We defer it to Coro. 4.5 after presenting our main results. for existing online RLHF algorithms (Xiong et al., 2024; Xie et al., 2024; Cen et al., 2024; Zhang et al., 2024), but also reveals a fundamental difference from the pure reward maximization setting, where lower bounds depending on those structural complexity factors have been established (Auer et al., 2002; Dani et al., 2008). We defer detailed discussions to Appx. E.3.
@@ 4 Provably Efficient Transfer Learning
Outline of Main Algorithm  Our main algorithm TPO is provided in Alg. 1, which leverages Alg. 2Transfer Policy Selection (TPS)—as a subroutine to select source policies to transfer from. TPO can be regarded as a mixture of standard online learning and transfer learning, balanced through a hyper-parameter $\alpha\in(0,1)$. Motivated by the implication of Thm. 3.2, TPO returns the detailed policy computed with all the data collected. For convenience, we divide the total number of iterations $T$ into $K=T/N$ blocks, each containing $N$ sub-iterations. In each block, we first run $\alpha N$ iterations of an OnLine learning algorithm $\texttt{Alg}_{\texttt{OL}}$, followed by $(1-\alpha)N$ iterations of transfer learning with policy selected by Alg. 2. Here, $\texttt{Alg}_{\texttt{OL}}$ can be any online algorithm with per-step no-regret guarantees, for example, XPO in Lem. 2.4. To save space, we defer to Appx. D the formal behavior assumption on $\texttt{Alg}_{\texttt{OL}}$ (Def. D.2) and concrete examples with verifications.
A Preview of Main Theorem  Before diving into the details, we first highlight the benefits of transfer learning by presenting an informal corollary regarding the regret bound of TPO, under concrete choices of $\texttt{Alg}_{\texttt{OL}}$ and $\alpha$.
Corollary 4.1.
Choosing XPO (Xie et al., 2024) as $\texttt{Alg}_{\texttt{OL}}$ and $\alpha=e^{-\frac{R}{\beta}}$, TPO achieves $\widetilde{O}(W\sqrt{T})$ regret when $T$ is small and $\widetilde{O}(\sqrt{T})$ regret after $T$ is large enough.
Coro. 4.1 is implied by our main result in Thm. 4.4. We refer to Remark F.3 for a detailed quantification of “small” and “large enough”. Compared with previous online RLHF results without transfer learning, in the early stage, TPO improves the structural complexity measure coefficients (e.g. ${\texttt{Cov}}_{\infty}$ in XPO) to the number of source tasks $W$, which is usually much smaller. Besides, it even gets rid of such coefficient term and achieves $\widetilde{O}(\sqrt{T})$ regret over time.
Next, we take a closer look at the transfer policy selection steps in Alg. 2 in Sec. 4.1, and provide detailed analyses and discussion of TPO in Sec. 4.2.
@@
             if $n\leq\alpha N$ then  $\pi^{k,n}\leftarrow\texttt{Alg}_{\texttt{OL}}(\alpha T,\Pi,\delta;\mathcal{D}^{k,n}_{\texttt{OL}})$  else   $\pi^{k,n}\leftarrow{\text{TPS}}(T,\Pi,\delta,\{r^{w}\}_{w=1}^{W};\mathcal{D}^{k,n})$
@@
We first clarify some notation. Given a dataset $\mathcal{D}:=\{(s^{i},a^{i},\widetilde{a}^{i},y^{i},\pi^{i})\}_{i=1}^{|\mathcal{D}|}$, $L_{\mathcal{D}}(r)$ denotes the average negative log-likelihood (NLL) loss regarding the reward $r$:
@@
$\pi_{\text{Dstl}},{\widehat{r}}_{\text{Dstl}}\leftarrow{\text{RPO}}(\text{conv}(\Pi),\mathcal{R}^{\Pi},\mathcal{D},\eta)$ with $\eta=c\cdot(1+e^{{R}})^{-2}\sqrt{\frac{1}{|\mathcal{D}|}\log\frac{|\Pi|T}{\delta}}.$
@@
4.2 Main Theorem and Interpretation
@@
We decompose the total regret into two parts depending on their origins. $\text{Reg}_{\texttt{OL}}^{(t)}$ comes from the regret by running the online algorithm $\texttt{Alg}_{\texttt{OL}}$. It is weighted by $\alpha$ since we only allocate $\alpha$-proportion of the samples for $\texttt{Alg}_{\texttt{OL}}$. $\text{Reg}_{\texttt{Trf}}^{(t)}$ represents the regret from transfer policies. The first term in Eq. (8) reflects the benefits of utilizing transfer policy over online learning. Here $\Delta_{\min}$ is contributed by source reward models $\{r^{w}\}_{w\in[W]}$, and the term $\iota^{k(\tau),n(\tau)}$ is due to the “self-transfer policy” $\pi_{\text{Dstl}}$, as we derived in Lem. 4.3. The second term in Eq. (8) results from the imperfection of source reward models: without prior knowledge on their quality, additional cost has to be paid during exploration.
@@
Stage 1: $t<\frac{W^{2}}{\Delta^{2}_{\min}}$  This corresponds to the early learning stage, when $t$ is relatively small. In this case, Thm. 4.4 implies the following regret bound:
@@
Input: $K$, $N$ and $\{r^{w}\}_{w\in[W]}$; Initialize: $\pi^{1}_{\texttt{OL}}\leftarrow\pi_{\texttt{ref}}$;
@@
             $\forall w\in[W],~{}\widehat{\text{WR}}^{\pi^{*}_{r^{w}}}\leftarrow\widehat{{\mathbb{P}}}_{r^{*}}^{k,n}(\pi^{*}_{r^{w}}\succ\pi^{k}_{\texttt{OL}})+c\sqrt{\log\frac{1}{\delta}/N^{k,n}(\pi^{*}_{r^{w}})}$;
@@
Remarks on choice of $\alpha$  We treat $\alpha$ as a hyperparameter. Without prior knowledge, a simple choice is $\alpha=e^{-\frac{R}{\beta}}$ with guarantee in Coro. 4.1. Under prior beliefs that high-quality source reward models are available, we may prefer smaller $\alpha$. Besides, due to the self-transfer learning, it is wise to gradually decay $\alpha$ to 0 as the iteration number grows.
Improved regret bound for standard online RLHF as a implication Although our focus is transfer learning, the standard online RLHF setting can be recovered when no source tasks are present, i.e., $W=0$. As stated below, TPO achieves $\widetilde{O}(\sqrt{T})$ regret over time by purely utilizing “self-transfer learning”, thereby strictly improving existing results (Xiong et al., 2024; Xie et al., 2024; Cen et al., 2024; Zhang et al., 2024) [^4] While omitted in our result, the dependence on the log-covering number (e.g., $\log|\Pi|$) matches with those previous works..
Corollary 4.5.
When $W=0$, under the same condition of Thm. 4.4, TPO reduces to a standard online RLHF method with $\widetilde{O}(\sqrt{T})$ regret after finite time.
@@ 5 From Theory to an Empirical Algorithm
Under BT-model555 Lem. 5.1 can be generalized beyond BT-model. Besides, it is possible to construct a lower bound involving ${\mathbb{P}}_{r^{*}}(\bar{\pi}\succ\pi)$ instead. See Lem. G.3 and Remark G.4 in Appx. G for more details. , for any $\pi$: ${\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi}\geq\!$$\max_{\gamma>0,\bar{\pi}}$$({\sqrt{(\gamma\!+\!2{\mathbb{P}}_{r^{*}}(\pi\!\succ\!\bar{\pi}))\log\frac{1+\gamma}{\gamma}}+\sqrt{\frac{J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\bar{\pi})}{2\beta}}})^{-1}.$
@@ 6 Experiments
Experiment Results and Discussion  We run Alg. 3 for $K=3$ iterations and compare its performance with three baselines: (I) vanilla iterative-DPO without transfer learning (i.e., setting $\text{Alg}_{\text{PO}}$ = DPO and $W=0$); (II) purely exploiting the worst source reward—ROUGE score; (III) purely exploiting the best source reward—T5-large. Concretely, baseline (I) removes the transfer learning component in Alg. 3 by assigning $\pi^{k,n}=\pi_{\texttt{OL}}^{k}$ for all $n\in[N]$. For baselines (II) and (III), the worst (ROUGE-LSum) and best (T5-Large) reward models from the candidates (a)-(d) are selected, and pure transfer learning is then performed using the responses recommended by the chosen reward models, i.e., $\pi^{k,n}=\pi^{*}_{r^{w}}$ in Alg. 3 for the selected $r^{w}$. Here the worst and best reward models are selected based on the final policy value when aligning with the given reward model.
@@ Appendix C Offline Learning Results in Previous Literature
${\text{RPO}}(\widetilde{\Pi},\mathcal{R},\mathcal{D},\eta)=\arg\max_{\pi\in\widetilde{\Pi}}\min_{r\in\mathcal{R}}L_{\mathcal{D}}(r)+\eta\mathbb{E}_{s\sim\rho,a\sim\pi,\widetilde{a}\sim\pi_{\texttt{ref}}}[r(s,a)-r(s,\widetilde{a})]-\beta\text{KL}(\pi\|\pi_{\texttt{ref}}),$ (12)
@@ F.2 Some Useful Lemmas
F.3 Proof for Thm. 4.4
See 4.4 Throught the proof, we follow the convention that $1/0=+\infty$.
@@
We first bound the accumulative error when $\mathbb{I}[\mathcal{E}^{k,n}_{\text{Dstl}}]=1$. On the good events in Lem. 4.2 and Lem. 4.3 (which holds w.p. $1-\delta$), $\mathbb{I}[\mathcal{E}^{k,n}_{\text{Dstl}}]=1$ implies
@@
Here we only focus on those source tasks with $\Delta(w)>0$, since transferring from $\pi^{*}_{r^{w}}$ with $\Delta(w)=0$ does not incur regret. We separate source tasks into two sets $\mathcal{W}_{\leq 2\Delta_{\min}}:=\{w\in[W]|\Delta(w)\leq 2\Delta_{\min}\}$ and $\mathcal{W}_{>2\Delta_{\min}}:=\{w\in[W]|\Delta(w)>2\Delta_{\min}\}$. For $w\in\mathcal{W}_{>2\Delta_{\min}}$, on the same good events in Lem. 4.2 and Lem. 4.3, $\mathbb{I}[\mathcal{E}^{n,k}_{w}]=1$ implies
@@
On the other hand, on the good events of Lem. 4.2 and Lem. 4.3, when ${\mathbb{I}}[\mathcal{E}^{n,k}_{w}\cap\mathcal{E}^{n,k}_{2\iota<\Delta_{\min}}]=1$, we must have:
$J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{ref}})-\iota^{k,n}\leq$ ${\widehat{V}}^{k,n}(\pi^{k,n}_{\text{Dstl}};\mathcal{D}^{k,n-1}{})$ (Lem. 4.2 and Lem. 4.3)
@@
$=$ $\widetilde{O}\Big{(}\sum_{(k,n)\in\mathcal{I}^{\texttt{Trf}}_{\widetilde{K},\widetilde{N}}}\Delta_{\min}\wedge\iota^{n,k}+{R}e^{2{R}}\sqrt{\alpha\mathcal{C}(\Pi)t_{\widetilde{K},\widetilde{N}}}+e^{2_{\max}}\sqrt{(1-\alpha)Wt_{\widetilde{K},\widetilde{N}}}\wedge\sum_{\begin{subarray}{c}w\in[W]\\ \Delta(w)>0\end{subarray}}\frac{e^{4{R}}}{\Delta(w)}\Big{)},$
@@
Remark F.3.
Under the choices of $\texttt{Alg}_{\texttt{OL}}\leftarrow$ XPO (Xie et al., 2024) and $\alpha=e^{-\frac{R}{\beta}}$, according to Eq. (9), TPO achieves $\widetilde{O}(W\sqrt{T})$ regret if $T<\frac{W^{2}}{\Delta^{2}_{\min}}$. On the other hand, by Lem. 3.1,
${\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi_{{\text{mix}}}^{t}}=1+\widetilde{O}(\kappa(e^{\frac{2R}{\beta}})\cdot\frac{\text{Reg}_{\texttt{OL}}^{(t)}+\text{Reg}_{\texttt{Trf}}^{(t)}}{\beta t})=1+\widetilde{O}(\kappa(e^{\frac{2R}{\beta}})\cdot\frac{1}{\alpha\beta}\sqrt{\frac{\mathcal{C}(\Pi)}{t}}).$
Therefore, consider $T_{0}=\widetilde{\Theta}(\frac{\kappa^{2}(e^{\frac{2R}{\beta}}\mathcal{C}(\Pi)}{\alpha^{2}\beta^{2}})$, where $\widetilde{\Theta}(\cdot)$ hides at most poly-log of $T$. At least when $T>T_{0}$, Eq. (10) implies $\text{Reg}_{\texttt{Trf}}^{(t)}=1+O(1)<C$ for some constant $C$. Since $\alpha=e^{-\frac{R}{\beta}}$, we also have $\text{Reg}_{\texttt{OL}}^{(T)}=\widetilde{O}(\sqrt{T})$ for any $T>0$. As a result, the total regret of TPO grows as $\widetilde{O}(\sqrt{T})$ as long as $T>T_{0}$.